# All genus open mirror symmetry for the projective line

### v1

@@ 1.1.2. Open Gromov-Witten theory for the projective line
The open Gromov-Witten invariants are more difficult to study than the closed case. The all genus $S^{1}$-equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ is studied by Buryak-Netser Zernik-Pandharipande-Tessler in [6]. A virtual localization formula is conjectured and proved in the genus zero case. The localization formula is crucial for the computation of open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. It is natural to ask whether one can extend the result in [19] to the open string sector. This will be introduced in Section 1.1.3.
@@ 1.1.3. All genus open mirror symmetry for the projective line
In this paper, we prove the all genus open mirror symmetry for $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$, extending the result in [19] to the open string sector. On A-model side, we consider the generating function of higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. On B-model side, instead of taking the oscillatory integrals, we take the expansion of the Chekhov-Eynard-Orantin invariants under certain local coordinate at the puncture of the mirror curve. The main theorem (Theorem 1.1) of this paper states that this expansion coincides with the generating function of higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ under the mirror map.
@@
We would like to remark that similar open mirror symmetry phenomenon appears in the case of toric Calabi-Yau 3-folds/3-orbifolds. Based on the work of Eynard-Orantin [15] and Mariño [33], Bouchard-Klemm-Mariño-Pasquetti (BKMP) [4, 5] proposed a new formalism of the topological B-model in terms of the Chekhov-Eynard-Orantin invariants of the mirror curve. BKMP conjectured a precise correspondence, known as the BKMP Remodeling Conjecture, between the local expansion of the Chekhov-Eynard-Orantin invariants at the puncture of the mirror curve and the generating function of open Gromov-Witten invariants of toric Calabi-Yau 3-folds/3-orbifolds. The mathematical study of the Remodeling Conjecture can be found, for example, in [3, 10, 16, 17, 18, 20, 21, 36, 37, 38, 39]. The difference in our case is that our target space $\mathbb{P}^{1}$ is compact and non-Calabi-Yau.
@@
We emphasize the following feature of our main result. Since $\mathbb{P}^{1}$ and $\mathbb{R}\mathbb{P}^{1}$ are compact, the non-equivariant limit of the $S^{1}$-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ is conjectured [6, Conjecture 1] to be equal to the non-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. These non-equivariant open Gromov-Witten invariants are conjectured to be defined geometrically via coherent boundary condition. In general, the computation of higher genus open Gromov-Witten invariants is quite difficult. The main result of our paper provides an effective algorithm of computing higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ recursively.
@@ 1.2. Statement of the main result
In Section 3.5, we will define the generating function $F_{g,h}(\mathbf{t},Q;X_{1},\dots,X_{h})$ of genus $g$, $h$ boundary circles open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. Here $\mathbf{t}=t^{0}1+t^{1}H$, $Q$ is the Novikov variable encoding the degree of the stable maps to $\mathbb{P}^{1}$, and $X_{1},\cdots,X_{h}$ are variables encoding the winding numbers (viewed as the open string coordinates).
@@
In Section 4.3, we will study the Chekhov-Eynard-Orantin invariants $\omega_{g,n}$ of the mirror curve $C_{q}$. Then we use $\omega_{g,n}$ to define the B-model potential $W_{g,n}(t^{0},q;X_{1},\dots,X_{n})$.
The following theorem is the main result of this paper:
@@ 1.3. Overview of the paper
In Section 2, we study the equivariant closed Gromov-Witten theory of $\mathbb{P}^{1}$ and give the graph sum formula for the descendant Gromov-Witten potential. In Section 3, we study the equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ and define the A-model potential $F_{g,n}$. In Section 4, we study the mirror curve of $\mathbb{P}^{1}$ and the Chekhov-Eynard-Orantin topological recursion. Then we use the Chekhov-Eynard-Orantin invariants $\omega_{g,n}$ to define the B-model potential $W_{g,n}$. In Section 5, we prove the all genus open mirror symmetry for $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$, which is the main result of this paper.
@@ Acknowledgements
The authors would like to thank Bohan Fang, Chiu-Chu Melissa Liu, and Song Yu for useful discussions. The second author is partially supported by the Natural Science Foundation of Beijing, China grant No. 1252008.
@@ 2.2. Equivariant Gromov-Witten invariants of $\mathbb{P}^{1}$
Let $E(\mathbb{P}^{1})$ denote the effective curve classes in $H_{2}(\mathbb{P}^{1};\mathbb{Z})$. Given nonnegative integers $g,n$ and an effective curve class $\beta\in E(\mathbb{P}^{1})$, let ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ be the moduli stack of genus $g$, $n$-pointed, degree $\beta$ stable maps to $\mathbb{P}^{1}$. Let $\mathrm{ev}_{i}:\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)\rightarrow\mathbb{P}^{1}$ be the evaluation map at the $i$-th marked point. The $T$-action on $\mathbb{P}^{1}$ induces an $T$-action on ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ and the evaluation map $\mathrm{ev}_{i}$ is $T$-equivariant.
For $i=1,\dots,n$, let $\mathbb{L}_{i}$ be the $i$-th tautological line bundle over ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ formed by the cotangent line at the $i$-th marked point. Define the $i$-th descendant class $\psi_{i}$ as
@@
$\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n}}(\gamma_{n})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}:=\sum_{\beta\in E(\mathbb{P}^{1})}\sum_{m=0}^{\infty}\frac{Q^{\beta}}{m!}\langle\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n}}(\gamma_{n}),\mathbf{t}^{m}\rangle_{g,n+m,\beta}^{\mathbb{P}^{1},T}.⟨$
@@
${\bf u}_{1},\dots,{\bf u}_{n}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}{\bf u}_{1}(\psi),\dots,{\bf u}_{n}(\psi)\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{a_{1},\dots,a_{n}\geq 0}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}(u_{1})_{a_{1}}\psi^{a_{1}},\dots,(u_{n})_{a_{n}}\psi^{a_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}.⟨$
@@
$\frac{\gamma_{1}}{z_{1}-\psi},\dots,\frac{\gamma_{n}}{z_{n}-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{a_{1},\dots,a_{n}\in\mathbb{Z}_{\geq 0}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\gamma_{1}\psi^{a_{1}},\dots,\gamma_{n}\psi^{a_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{i=1}^{n}z_{i}^{-a_{i}-1}.⟨$

### v2

@@ 1.1.2. Open Gromov-Witten theory for the projective line
The open Gromov-Witten invariants are more difficult to study than the closed case. The all genus $S^{1}$-equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ is studied via coherent boundary conditions by Buryak-Netser Zernik-Pandharipande-Tessler in [6]. A localization formula is conjectured and proved in the genus zero case. The localization formula is crucial for the computation of open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. It is natural to ask whether one can extend the result in [19] to the open string sector. This will be introduced in Section 1.1.3.
@@ 1.1.3. All genus open mirror symmetry for the projective line
In this paper, we prove the all genus open mirror symmetry for $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$, extending the result in [19] to the open string sector. On A-model side, we will study the open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ from two different but related points of view. From the first point of view, we define the higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via integration over the fixed locus of the $S^{1}$-action and consider the corresponding generating functions in Section 3. On B-model side, instead of taking the oscillatory integrals, we take the expansion of the Chekhov-Eynard-Orantin invariants under certain local coordinate at the puncture of the mirror curve. The first main theorem (Theorem 1.1) of this paper states that this expansion coincides with the generating function of higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ under the mirror map. The second main theorem (Theorem 1.2) generalizes the above result to the case when descendant insertions appear.
From the second point of view, we define the higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ as a graph sum in Section 6.2 and Section 6.3. The vertex factor in this graph sum formula is given by the open Gromov-Witten invariants via integration over the fixed locus defined in Section 3. In other words, the open Gromov-Witten invariants via integration over the fixed locus are the building blocks in the second point of view of the open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. In [6], this graph sum formula is conjectured to have a geometric definition via coherent boundary conditions. So we call the second point of view the open Gromov-Witten invariants via coherent boundary conditions. The third main theorem (Theorem 1.3) of this paper is the all genus mirror symmetry for open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via coherent boundary conditions.
@@
We would like to remark that similar open mirror symmetry phenomenon appears in the case of toric Calabi-Yau 3-folds/3-orbifolds, where the open Gromov-Witten invariants are defined via integration over the fixed locus. Based on the work of Eynard-Orantin [15] and Mariño [33], Bouchard-Klemm-Mariño-Pasquetti (BKMP) [4, 5] proposed a new formalism of the topological B-model in terms of the Chekhov-Eynard-Orantin invariants of the mirror curve. BKMP conjectured a precise correspondence, known as the BKMP Remodeling Conjecture, between the local expansion of the Chekhov-Eynard-Orantin invariants at the puncture of the mirror curve and the generating function of open Gromov-Witten invariants of toric Calabi-Yau 3-folds/3-orbifolds. The mathematical study of the Remodeling Conjecture can be found, for example, in [3, 10, 16, 17, 18, 20, 21, 36, 37, 38, 39]. The difference in our case is that our target space $\mathbb{P}^{1}$ is compact and non-Calabi-Yau.
@@
We emphasize the following feature of our main results. Since $\mathbb{P}^{1}$ and $\mathbb{R}\mathbb{P}^{1}$ are compact, the non-equivariant limit of the $S^{1}$-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ defined in Section 6 is conjectured [6, Conjecture 1] to be equal to the non-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$, defined geometrically via coherent boundary conditions. In general, the computation of higher genus open Gromov-Witten invariants is quite difficult. The main results of our paper provide an effective algorithm of computing higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ recursively.
@@ 1.2. Statement of the main result
In Section 3.5, we will define the generating function $F_{g,n}(\mathbf{t},Q;X_{1},\dots,X_{n})$ of genus $g$, $n$ boundary circles open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via integration over the fixed locus of the $S^{1}$-action. Here $\mathbf{t}=t^{0}1+t^{1}H$, $Q$ is the Novikov variable encoding the degree of the stable maps to $\mathbb{P}^{1}$, and $X_{1},\cdots,X_{n}$ are variables encoding the winding numbers (viewed as the open string coordinates). We will also define the generating function $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\cdots\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}(\mathbf{t},X_{1},\dots,X_{n})$ of genus $g$, $m$ descendant insertion classes, $n$ boundary circles open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$.
@@
In Section 4.3, we will study the Chekhov-Eynard-Orantin invariants $\omega_{g,n}$ of the mirror curve $C_{q}$. Then we use $\omega_{g,n}$ to define the B-model open primary potential $W_{g,n}(t^{0},q;X_{1},\dots,X_{n})$. In Section 4.4, we study the oscillatory integrals of $\omega_{g,n}$ and define the B-model open descendant potential $W_{g,m,n}(t^{0},q;\mathcal{E}_{1},\dots,\mathcal{E}_{m};X_{1},\dots,X_{n})$, where $\mathcal{E}_{1},\dots,\mathcal{E}_{m}\in K_{T}(\mathbb{P}^{1})$.
The following three theorems are the main results of this paper:
@@
Theorem 1.2 (= Theorem 5.8).
Suppose $g,m,n\geq 0$, for any $\mathcal{E}_{1},\dots,\mathcal{E}_{m}\in K_{T}(\mathbb{P}^{1})$, we have
$W_{g,m,n}(t^{0},q;\mathcal{E}_{1},\dots,\mathcal{E}_{m};X_{1},\dots,X_{n})$
$=(-1)^{g-1}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}},\dots,\frac{\kappa_{z_{m}}(\mathcal{E}_{m})}{z_{m}-\psi_{m}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}(\mathbf{t},X_{1},\dots,X_{n}).$
Here the characteristic class $\kappa_{z}$ is defined in Section 4.4 and $z_{1},\cdots,z_{m}$ are formal variables. The dependence of $W_{g,m,n}$ on $z_{1},\cdots,z_{m}$ is via oscillatory integrals defined in Section $4.$
In Section 6.1 and 6.2, we will study the moduli specification $S$ and the combinatorial graph sum formula of open Gromov-Witten invariants $\text{OGW}(S,\vec{a},\vec{\gamma})$, where the vertex factor is given by the open Gromov-Witten invariants via integration over the fixed locus defined in Section 3. In [6], $\text{OGW}(S,\vec{a},\vec{\gamma})$ is conjectured to have a geometric definition via coherent boundary conditions. We will define the decorated graph $\Gamma_{S}$ and the generating function $\text{OGW}(\Gamma_{S},\vec{a},\vec{\gamma},\mathbf{t})$ in Section 6.3. In Section 6.4, we define $\text{OGW}_{g,m,n}(\Gamma_{S},\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}},\mathbf{t})$ as a function of $z_{i}$ composed of $\text{OGW}(\Gamma_{S},\vec{a},\vec{\gamma},\mathbf{t})$. The B-model potential $G_{g,m,n}(\Gamma_{S},\mathbf{t})$ is defined in Section 6.4 via graph sum formula, where the building blocks are Chekhov-Eynard-Orantin invariants. The third main theorem is the mirror symmetry for $\text{OGW}_{g,m,n}(\Gamma_{S},\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}},\mathbf{t})$.
Theorem 1.3 (= Theorem 6.9).
Let $S$ be a pure connected moduli specification with one black vertex, $n$ white vertices and $m$ marked labels $\sf I$}. Then we have
$\text{OGW}_{g,m,n}(\Gamma_{S},\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}},\mathbf{t})=G_{g,m,n}(\Gamma_{S},\mathbf{t}).$
@@ 1.3. Overview of the paper
In Section 2, we study the equivariant closed Gromov-Witten theory of $\mathbb{P}^{1}$ and give the graph sum formula for the descendant Gromov-Witten potential. In Section 3, we define the equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via integration over the fixed locus of the $S^{1}$-action and define the A-model potential $F_{g,n}$ and $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}},\dots,\frac{\kappa_{z_{m}}(\mathcal{E}_{m})}{z_{m}-\psi_{m}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}$. In Section 4, we study the mirror curve of $\mathbb{P}^{1}$ and the Chekhov-Eynard-Orantin topological recursion. Then we use the Chekhov-Eynard-Orantin invariants $\omega_{g,n}$ to define the B-model potential $W_{g,m,n}$. In Section 5, we prove the all genus open mirror symmetry for $F_{g,n}$ and $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}},\dots,\frac{\kappa_{z_{m}}(\mathcal{E}_{m})}{z_{m}-\psi_{m}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}$. In Section 6, we consider the open Gromov-Witten invariants $\text{OGW}(S,\vec{a},\vec{\gamma})$ and prove the corresponding all genus open mirror symmetry.
@@ Acknowledgements
The authors would like to thank Bohan Fang, Chiu-Chu Melissa Liu, and Song Yu for useful discussions. The second author is partially supported by the Natural Science Foundation of Beijing, China (grant No. 1252008) and NSFC (grant No. 12571067).
@@ 2.2. Equivariant Gromov-Witten invariants of $\mathbb{P}^{1}$
Let $E(\mathbb{P}^{1})$ denote the effective curve classes in $H_{2}(\mathbb{P}^{1};\mathbb{Z})$. Given nonnegative integers $g,n$ and an effective curve class $\beta\in E(\mathbb{P}^{1})$, let $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ be the moduli stack of genus $g$, $n$-pointed, degree $\beta$ stable maps to $\mathbb{P}^{1}$. Let $\mathrm{ev}_{i}:\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)\rightarrow\mathbb{P}^{1}$ be the evaluation map at the $i$-th marked point. The $T$-action on $\mathbb{P}^{1}$ induces an $T$-action on $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ and the evaluation map $\mathrm{ev}_{i}$ is $T$-equivariant.
For $i=1,\dots,n$, let $\mathbb{L}_{i}$ be the $i$-th tautological line bundle over $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ formed by the cotangent line at the $i$-th marked point. Define the $i$-th descendant class $\psi_{i}$ as
@@
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n}}(\gamma_{n})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}:=\sum_{\beta\in E(\mathbb{P}^{1})}\sum_{m=0}^{\infty}\frac{Q^{\beta}}{m!}\langle\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n}}(\gamma_{n}),\mathbf{t}^{m}\rangle_{g,n+m,\beta}^{\mathbb{P}^{1},T}.$
@@
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}{\bf u}_{1},\dots,{\bf u}_{n}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}{\bf u}_{1}(\psi),\dots,{\bf u}_{n}(\psi)\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{a_{1},\dots,a_{n}\geq 0}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}(u_{1})_{a_{1}}\psi^{a_{1}},\dots,(u_{n})_{a_{n}}\psi^{a_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}.$
@@
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\gamma_{1}}{z_{1}-\psi},\dots,\frac{\gamma_{n}}{z_{n}-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{a_{1},\dots,a_{n}\in\mathbb{Z}_{\geq 0}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\gamma_{1}\psi^{a_{1}},\dots,\gamma_{n}\psi^{a_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{i=1}^{n}z_{i}^{-a_{i}-1}.$